OpenStudy (anonymous):

Find the value of 32 yd^3 in ft^3

OpenStudy (anonymous):

32* 27=864 cu ft

OpenStudy (wolf1728):

1 yard = 3 feet 1 cubic yard = 3 * 3 * 3 = 27 cubic feet 32 cubic yards = 32*27 cubic feet =864 cubic feet

OpenStudy (anonymous):

Thanks so much !

OpenStudy (anonymous):

I have another one that's asking find the value of 294cm^2 in m^2

OpenStudy (anonymous):

@robtobey @wolf1728

OpenStudy (anonymous):

\[\frac{294}{100^2}=0.0294 \text{ sq} \text{ meters}\]

OpenStudy (anonymous):

thanks again your the best(: